# Parabolic induction in the homotopy category of pro-$p$ Iwahori-Hecke modules

### v2

@@ 4.1. Constructing model structures from a recollement
    If the model structure on ${\mathcal{A}}$ is injective, then ${\mathcal{A}}_{i}^{\mathcal{S}}=({\mathcal{W}}_{\mathcal{S}},{\mathcal{F}}_{\mathcal{A}},{\mathcal{F}}_{\mathcal{S}})$ defines a hereditary, cofibrantly generated, abelian model structure on ${\mathcal{A}}$ whose homotopy category is canonically equivalent to ${\mathcal{S}}$. Furthermore, the adjunction ${\operatorname{id}}:{\mathcal{A}}_{i}^{\mathcal{S}}\rightleftarrows{\mathcal{A}}:{\operatorname{id}}$ is Quillen and its total derived adjunction is naturally isomorphic to the adjoint pair $(i,\mathbf{R}{\mathsf{R}})$.
@@
    If the model structure on ${\mathcal{A}}$ is projective, then ${\mathcal{A}}_{p}^{\mathcal{S}}=({\mathcal{C}}_{\mathcal{A}},{\mathcal{W}}_{\mathcal{S}},{\mathcal{C}}_{\mathcal{S}})$ defines a hereditary, cofibrantly generated, abelian model structure on ${\mathcal{A}}$ whose homotopy category is canonically equivalent to ${\mathcal{S}}$. Furthermore, the adjunction ${\operatorname{id}}:{\mathcal{A}}\rightleftarrows{\mathcal{A}}_{p}^{\mathcal{S}}:{\operatorname{id}}$ is Quillen and its total derived adjunction is naturally isomorphic to the adjoint pair $(\mathbf{L}{\mathsf{L}},i)$.
@@
Everything follow immediately from 4.2 and [7, Propositions 1.4.2 & 1.4.6], except the cofibrant generation for which we need to check that the associated cotorsion pairs are small. In the injective case we have $({\mathcal{W}}_{\mathcal{S}}\cap{\mathcal{F}}_{\mathcal{S}},{\mathcal{F}}_{\mathcal{A}})=({\mathcal{W}}_{\mathcal{A}},{\mathcal{F}}_{\mathcal{A}})$ and in the projective case we have $({\mathcal{C}}_{\mathcal{A}},{\mathcal{W}}_{\mathcal{S}}\cap{\mathcal{C}}_{\mathcal{S}})=({\mathcal{C}}_{\mathcal{A}},{\mathcal{W}}_{\mathcal{A}})$ , c.f. 4.5, and so it follows from that these are small. Furthermore, the proof 4.2 shows that $({\mathcal{W}}_{\mathcal{S}},{\mathcal{F}}_{\mathcal{A}}\cap{\mathcal{F}}_{\mathcal{S}})$, in the injective case, and $({\mathcal{C}}_{\mathcal{A}}\cap{\mathcal{C}}_{\mathcal{S}},{\mathcal{W}}_{\mathcal{S}})$, in the projective case, are small as well. ∎
@@
Suppose that ${\mathcal{A}}$ is a Grothendieck abelian category equipped with both an injective model structure $({\mathcal{A}},{\mathcal{F}}_{\mathcal{A}},W_{\mathcal{A}})$ and a projective model structure $({\mathcal{C}}_{\mathcal{A}},{\mathcal{A}},{\mathcal{W}}_{\mathcal{A}})$ with the same classes of trivial objects, and let ${\operatorname{Ho}}({\mathcal{A}})$ denote their common homotopy category. Assume that ${\mathcal{S}}$ is a thick subcategory of ${\operatorname{Ho}}({\mathcal{A}})$ satisfying the requirements of 4.1. Let ${\mathcal{W}}_{\mathcal{S}}$, ${\mathcal{F}}_{\mathcal{S}}$ and ${\mathcal{C}}_{\mathcal{S}}$ be as in 4.2.
@@
    ${{}^{\ddagger}}{\mathcal{A}}^{i}_{\mathcal{S}}=({\mathcal{A}},{\mathcal{W}}_{\mathcal{S}}\cap{\mathcal{F}}_{\mathcal{A}},{\mathcal{C}}_{\mathcal{S}})$ defines an injective, cofibrantly generated, abelian model structure on ${\mathcal{A}}$. Furthermore, the identity ${\operatorname{id}}:{{}^{\ddagger}}{\mathcal{A}}^{i}_{\mathcal{S}}\to{\mathcal{A}}_{i}^{\mathcal{S}}$ is a right Quillen equivalence, the adjunction ${\operatorname{id}}:{\mathcal{A}}\rightleftarrows{{}^{\ddagger}}{\mathcal{A}}^{i}_{\mathcal{S}}:{\operatorname{id}}$ is Quillen (for the injective model structure on ${\mathcal{A}}$) and its total derived adjunction is naturally isomorphic to the adjoint pair $(\mathbf{L}{\mathsf{L}},i)$.
@@
    ${{}^{\ddagger}}{\mathcal{A}}^{p}_{\mathcal{S}}=({\mathcal{W}}_{\mathcal{S}}\cap{\mathcal{C}}_{\mathcal{A}},{\mathcal{A}},{\mathcal{F}}_{\mathcal{S}})$ defines a projective, cofibrantly generated, abelian model structure on ${\mathcal{A}}$. Furthermore, the identity ${\operatorname{id}}:{{}^{\ddagger}}{\mathcal{A}}^{p}_{\mathcal{S}}\to{\mathcal{A}}_{p}^{\mathcal{S}}$ is a left Quillen equivalence, the adjunction ${\operatorname{id}}:{{}^{\ddagger}}{\mathcal{A}}^{p}_{\mathcal{S}}\rightleftarrows{\mathcal{A}}:{\operatorname{id}}$ is Quillen (for the projective model structure on ${\mathcal{A}}$) and its total derived adjunction is naturally isomorphic to the adjoint pair $(i,\mathbf{R}{\mathsf{R}})$.
@@
(i) By 4.5, the pair $({\mathcal{A}},{\mathcal{W}}_{\mathcal{S}}\cap{\mathcal{F}}_{\mathcal{A}}\cap{\mathcal{C}}_{\mathcal{S}})=({\mathcal{A}},\mathrm{Inj}({\mathcal{A}}))$ is a small cotorsion pair. Furthermore, by assumption and by 4.7(ii) we have that both ${\mathcal{W}}_{\mathcal{S}}$ an ${\mathcal{F}}_{\mathcal{A}}$ are of the form $\mathcal{G}^{\perp}$, hence so is their intersection. Thus we only need to show that ${\mathcal{C}}_{\mathcal{S}}={{}^{\perp}}({\mathcal{W}}_{\mathcal{S}}\cap{\mathcal{F}}_{\mathcal{A}})$ to get a small cotorsion pair. Since every object of ${\mathcal{W}}_{\mathcal{S}}$ is isomorphic in ${\operatorname{Ho}}({\mathcal{A}})$ to an object of of ${\mathcal{W}}_{\mathcal{S}}\cap{\mathcal{F}}_{\mathcal{A}}$ and since those are fibrant in ${\mathcal{A}}$, we may apply 2.16 to deduce that an object lies in ${{}^{\perp}}({\mathcal{W}}_{\mathcal{S}}\cap{\mathcal{F}}_{\mathcal{A}})$ if and only if its image in ${\operatorname{Ho}}({\mathcal{A}})$ lies in ${}^{\ddagger}{\mathcal{S}}={\operatorname{Ker}}(\mathbf{L}{\mathsf{L}})$ as required.
@@
(ii) Again, by 4.5 the pair $({\mathcal{W}}_{\mathcal{S}}\cap{\mathcal{C}}_{\mathcal{A}}\cap{\mathcal{F}}_{\mathcal{S}},{\mathcal{A}})=(\mathrm{Proj}({\mathcal{A}}),{\mathcal{A}})$ is a small cotorsion pair. Furthermore, by 4.7(i) there exists a set $\mathcal{G}\subseteq{\mathcal{W}}_{\mathcal{S}}$ such that $\mathcal{G}^{\perp}={\mathcal{F}}_{\mathcal{A}}\cap{\mathcal{F}}_{\mathcal{S}}$. Now define a set $\mathcal{G}^{\prime}\subseteq{\mathcal{W}}_{\mathcal{S}}\cap{\mathcal{C}}_{\mathcal{A}}$ by letting the objects of $\mathcal{G}^{\prime}$ be a choice of cofibrant replacements in ${\mathcal{A}}$ for the objects of $\mathcal{G}$. Since these are cofibrant, we may apply 2.16 again to deduce that $(\mathcal{G}^{\prime})^{\perp}={\mathcal{F}}_{\mathcal{S}}$. Furthermore, the pair $({{}^{\perp}}{\mathcal{F}}_{\mathcal{S}},{\mathcal{F}}_{\mathcal{S}})$ is now a small cotorsion pair, such that ${{}^{\perp}}{\mathcal{F}}_{\mathcal{S}}$ is the closure of $\mathcal{G}^{\prime}\cup\{Z\}$ under transfinite compositions and direct summands by [15, Lemma 6.9], where $Z$ is a choice of projective generator for ${\mathcal{A}}$. Since both ${\mathcal{W}}_{\mathcal{S}}$ and ${\mathcal{C}}_{\mathcal{A}}$ are closed under those by Lemma 6.2 in loc. cit. and 4.7(i), we have ${{}^{\perp}}{\mathcal{F}}_{\mathcal{S}}\subseteq{\mathcal{W}}_{\mathcal{S}}\cap{\mathcal{C}}_{\mathcal{A}}$. To show equality it therefore suffices to show that an object of ${\mathcal{C}}_{\mathcal{A}}$ lies in ${{}^{\perp}}{\mathcal{F}}_{\mathcal{S}}$ if and only if it lies in ${\mathcal{W}}_{\mathcal{S}}$. This is proved via the usual argument using 2.16. The statements on the Quillen functors are proved dually to (i).
@@
Finally, ${\operatorname{id}}:{{}^{\ddagger}}{\mathcal{A}}^{i}_{\mathcal{S}}\to{{}^{\ddagger}}{\mathcal{A}}^{p}_{\mathcal{S}}$ will by definition be a right Quillen equivalence if, for $X\in{\mathcal{W}}_{\mathcal{S}}\cap{\mathcal{C}}_{\mathcal{A}}$ and $Y\in{\mathcal{W}}_{\mathcal{S}}\cap{\mathcal{F}}_{\mathcal{A}}$, a map $f:X\to Y$ is a weak equivalence in ${{}^{\ddagger}}{\mathcal{A}}^{p}_{\mathcal{S}}$ if and only if it is a weak equivalence in ${{}^{\ddagger}}{\mathcal{A}}^{i}_{\mathcal{S}}$. Note that the trivial cofibrations in ${{}^{\ddagger}}{\mathcal{A}}^{p}_{\mathcal{S}}$ are precisely the monomorphisms with projective cokernel, and so are still weak equivalences in ${{}^{\ddagger}}{\mathcal{A}}^{i}_{\mathcal{S}}$. Thus, by factorising $f$ in ${{}^{\ddagger}}{\mathcal{A}}^{p}_{\mathcal{S}}$ as a trivial cofibration followed by a fibration we may assume that $f$ is a fibration. Then it is a weak equivalence in ${{}^{\ddagger}}{\mathcal{A}}^{p}_{\mathcal{S}}$ if and only if its kernel lies in ${\mathcal{F}}_{\mathcal{S}}$. But since both $X$ and $Y$ lie in ${\mathcal{W}}_{\mathcal{S}}$ by assumption and ${\mathcal{W}}_{\mathcal{S}}$ is thick, this happens if and only if $\ker(f)\in{\mathcal{W}}_{\mathcal{S}}\cap{\mathcal{F}}_{\mathcal{S}}={\mathcal{W}}_{\mathcal{A}}$, c.f. 4.5. Similarly, $f$ is a weak equivalence in ${{}^{\ddagger}}{\mathcal{A}}^{i}_{\mathcal{S}}$ if and only if $\ker(f)\in{\mathcal{W}}_{\mathcal{S}}\cap{\mathcal{C}}_{\mathcal{S}}={\mathcal{W}}_{\mathcal{A}}$. ∎
@@
In particular, by considering ${\mathcal{A}}={\operatorname{Mod}}({\mathcal{H}})$ with its Gorenstein projective and injective model structures and by letting =${\mathcal{M}}=\{M_{1},\ldots,M_{n}\}$ be some fixed collection standard Levi subgroups of $G$, the above tells us that the recollement in 3.13(i) may be reconstructed from various model structures on ${\operatorname{Mod}}({\mathcal{H}})$ by only using the Gorenstein projective model structure. If we define
@@
where $F_{M}({\mathfrak{m}}):={\mathbb{X}}_{M}\otimes_{{\mathcal{H}}_{M}}{\mathfrak{m}}$ and $U_{M}(V):={\operatorname{Hom}}_{M}({\mathbb{X}}_{M},V)\cong V^{I_{M}}$ for any ${\mathfrak{m}}\in{\operatorname{Mod}}({\mathcal{H}}_{M})$ and $V\in{\operatorname{Rep}}_{k}^{\infty}(M)$. When $M=G$, we drop the subscript and write $F$ and $U$ for the corresponding functors. On the smooth representations side, the parabolic induction functor ${\operatorname{Ind}}_{P}^{G}:{\operatorname{Rep}}_{k}^{\infty}(M)\to{\operatorname{Rep}}_{k}^{\infty}(G)$ is defined by first inflating smooth representations of $M\cong P/N$ to $P$ and then applying the smooth induction functor. This functor has both a left and right adjoint (c.f. [31, Proposition 4.3]), and its left adjoint can be described explicitly as the Jacquet functor $V\mapsto J_{P}^{G}(V):=V_{N}$ where $V_{N}$ denotes the $N$-coinvariants of $V$. This construction is compatible with the parabolic induction for ${\mathcal{H}}$-modules by a result of Ollivier-Vignéras.
Theorem 4.10 ([27, Theorem 1]).
@@
For any collection ${\mathcal{M}}$ standard Levi subgroups of $G$, the right transfer of the model structures ${\operatorname{Mod}}({\mathcal{H}})^{G}_{\mathcal{M}}$ and ${{}^{\ddagger}}{\operatorname{Mod}}({\mathcal{H}})^{G}_{\mathcal{M}}$ to ${\operatorname{Rep}}_{k}^{\infty}(G)$ along $U$ both exist.

### v4

@@ 4.1. Constructing model structures from a recollement
    If the model structure on ${\mathcal{A}}$ is injective, then ${\mathcal{A}}_{i}^{\mathcal{S}}=({\mathcal{W}}_{\mathcal{S}},{\mathcal{F}}_{\mathcal{A}},{\mathcal{F}}_{\mathcal{S}})$ defines a hereditary, cofibrantly generated, abelian model structure on ${\mathcal{A}}$ whose homotopy category is canonically equivalent to ${\mathcal{S}}$. Furthermore, the adjunction ${\operatorname{id}}:{\mathcal{A}}_{i}^{\mathcal{S}}\rightleftarrows{\mathcal{A}}:{\operatorname{id}}$ is Quillen and its total derived adjunction is naturally isomorphic to the adjoint pair $(i,{\mathsf{R}}_{{\mathcal{S}}})$.
@@
    If the model structure on ${\mathcal{A}}$ is projective, then ${\mathcal{A}}_{p}^{\mathcal{S}}=({\mathcal{C}}_{\mathcal{A}},{\mathcal{W}}_{\mathcal{S}},{\mathcal{C}}_{\mathcal{S}})$ defines a hereditary, cofibrantly generated, abelian model structure on ${\mathcal{A}}$ whose homotopy category is canonically equivalent to ${\mathcal{S}}$. Furthermore, the adjunction ${\operatorname{id}}:{\mathcal{A}}\rightleftarrows{\mathcal{A}}_{p}^{\mathcal{S}}:{\operatorname{id}}$ is Quillen and its total derived adjunction is naturally isomorphic to the adjoint pair $({\mathsf{L}}_{{\mathcal{S}}},i)$.
@@
Everything follows immediately from 4.2 and [7, Propositions 1.4.2 & 1.4.6], except the cofibrant generation for which we need to check that the associated cotorsion pairs are small. In the injective case we have $({\mathcal{W}}_{\mathcal{S}}\cap{\mathcal{F}}_{\mathcal{S}},{\mathcal{F}}_{\mathcal{A}})=({\mathcal{W}}_{\mathcal{A}},{\mathcal{F}}_{\mathcal{A}})$ and in the projective case we have $({\mathcal{C}}_{\mathcal{A}},{\mathcal{W}}_{\mathcal{S}}\cap{\mathcal{C}}_{\mathcal{S}})=({\mathcal{C}}_{\mathcal{A}},{\mathcal{W}}_{\mathcal{A}})$ , cf. 4.5, and so it follows from our assumptions that these are small. Furthermore, the proof of 4.2 shows that $({\mathcal{W}}_{\mathcal{S}},{\mathcal{F}}_{\mathcal{A}}\cap{\mathcal{F}}_{\mathcal{S}})$, in the injective case, and $({\mathcal{C}}_{\mathcal{A}}\cap{\mathcal{C}}_{\mathcal{S}},{\mathcal{W}}_{\mathcal{S}})$, in the projective case, are small as well. ∎
@@
Suppose that ${\mathcal{A}}$ is a Grothendieck abelian category equipped with both an injective model structure $({\mathcal{A}},{\mathcal{F}}_{\mathcal{A}},{\mathcal{W}}_{\mathcal{A}})$ and a projective model structure $({\mathcal{C}}_{\mathcal{A}},{\mathcal{A}},{\mathcal{W}}_{\mathcal{A}})$ with the same classes of trivial objects, and assume that these are both cofibrantly generated. Let ${\operatorname{Ho}}({\mathcal{A}})$ denote their common homotopy category and assume ${\mathcal{S}}$ is a thick subcategory of ${\operatorname{Ho}}({\mathcal{A}})$ satisfying the requirements of 4.1. Let ${\mathcal{W}}_{\mathcal{S}}$, ${\mathcal{F}}_{\mathcal{S}}$ and ${\mathcal{C}}_{\mathcal{S}}$ be as in 4.2.
@@
    ${{}^{\ddagger}}{\mathcal{A}}^{i}_{\mathcal{S}}=({\mathcal{A}},{\mathcal{W}}_{\mathcal{S}}\cap{\mathcal{F}}_{\mathcal{A}},{\mathcal{C}}_{\mathcal{S}})$ defines an injective, cofibrantly generated, abelian model structure on ${\mathcal{A}}$. Furthermore, the identity ${\operatorname{id}}:{{}^{\ddagger}}{\mathcal{A}}^{i}_{\mathcal{S}}\to{\mathcal{A}}_{i}^{\mathcal{S}}$ is a right Quillen equivalence, the adjunction ${\operatorname{id}}:{\mathcal{A}}\rightleftarrows{{}^{\ddagger}}{\mathcal{A}}^{i}_{\mathcal{S}}:{\operatorname{id}}$ is Quillen (for the injective model structure on ${\mathcal{A}}$) and its total derived adjunction is naturally isomorphic to the adjoint pair $({\mathsf{L}}_{{\mathcal{S}}},i)$.
@@
    ${{}^{\ddagger}}{\mathcal{A}}^{p}_{\mathcal{S}}=({\mathcal{W}}_{\mathcal{S}}\cap{\mathcal{C}}_{\mathcal{A}},{\mathcal{A}},{\mathcal{F}}_{\mathcal{S}})$ defines a projective, cofibrantly generated, abelian model structure on ${\mathcal{A}}$. Furthermore, the identity ${\operatorname{id}}:{{}^{\ddagger}}{\mathcal{A}}^{p}_{\mathcal{S}}\to{\mathcal{A}}_{p}^{\mathcal{S}}$ is a left Quillen equivalence, the adjunction ${\operatorname{id}}:{{}^{\ddagger}}{\mathcal{A}}^{p}_{\mathcal{S}}\rightleftarrows{\mathcal{A}}:{\operatorname{id}}$ is Quillen (for the projective model structure on ${\mathcal{A}}$) and its total derived adjunction is naturally isomorphic to the adjoint pair $(i,{\mathsf{R}}_{{\mathcal{S}}})$.
@@
(i) By 4.5, the pair $({\mathcal{A}},{\mathcal{W}}_{\mathcal{S}}\cap{\mathcal{F}}_{\mathcal{A}}\cap{\mathcal{C}}_{\mathcal{S}})=({\mathcal{A}},\mathrm{Inj}({\mathcal{A}}))$ is a small cotorsion pair. Furthermore, by assumption and by 4.7(i) we have that both ${\mathcal{W}}_{\mathcal{S}}$ and ${\mathcal{F}}_{\mathcal{A}}$ are of the form $\mathcal{G}^{\perp}$, hence so is their intersection. Thus we only need to show that ${\mathcal{C}}_{\mathcal{S}}={{}^{\perp}}({\mathcal{W}}_{\mathcal{S}}\cap{\mathcal{F}}_{\mathcal{A}})$ to get a small cotorsion pair. Since every object of ${\mathcal{W}}_{\mathcal{S}}$ is isomorphic in ${\operatorname{Ho}}({\mathcal{A}})$ to an object of ${\mathcal{W}}_{\mathcal{S}}\cap{\mathcal{F}}_{\mathcal{A}}$ and since those are fibrant in ${\mathcal{A}}$, we may apply 2.16 to deduce that an object lies in ${{}^{\perp}}({\mathcal{W}}_{\mathcal{S}}\cap{\mathcal{F}}_{\mathcal{A}})$ if and only if its image in ${\operatorname{Ho}}({\mathcal{A}})$ lies in ${}^{\ddagger}{\mathcal{S}}={\operatorname{Ker}}({\mathsf{L}}_{{\mathcal{S}}})$ as required.
@@
(ii) Again, by 4.5 the pair $({\mathcal{W}}_{\mathcal{S}}\cap{\mathcal{C}}_{\mathcal{A}}\cap{\mathcal{F}}_{\mathcal{S}},{\mathcal{A}})=(\mathrm{Proj}({\mathcal{A}}),{\mathcal{A}})$ is a small cotorsion pair. Furthermore, by 4.7(i) there exists a set $\mathcal{G}\subseteq{\mathcal{W}}_{\mathcal{S}}$ such that $\mathcal{G}^{\perp}={\mathcal{F}}_{\mathcal{A}}\cap{\mathcal{F}}_{\mathcal{S}}$. Now define a set $\mathcal{G}^{\prime}\subseteq{\mathcal{W}}_{\mathcal{S}}\cap{\mathcal{C}}_{\mathcal{A}}$ by letting the objects of $\mathcal{G}^{\prime}$ be a choice of cofibrant replacements in ${\mathcal{A}}$ for the objects of $\mathcal{G}$. Since these are cofibrant, we may apply 2.16 again to deduce that $(\mathcal{G}^{\prime})^{\perp}={\mathcal{F}}_{\mathcal{S}}$. Furthermore, the pair $({{}^{\perp}}{\mathcal{F}}_{\mathcal{S}},{\mathcal{F}}_{\mathcal{S}})$ is now a small cotorsion pair, such that ${{}^{\perp}}{\mathcal{F}}_{\mathcal{S}}$ is the closure of $\mathcal{G}^{\prime}\cup\{Z\}$ under transfinite compositions and direct summands by [17, Lemma 6.9], where $Z$ is a choice of projective generator for ${\mathcal{A}}$. Since both ${\mathcal{W}}_{\mathcal{S}}$ and ${\mathcal{C}}_{\mathcal{A}}$ are closed under those by Lemma 6.2 in loc. cit. and 4.7(i), we have ${{}^{\perp}}{\mathcal{F}}_{\mathcal{S}}\subseteq{\mathcal{W}}_{\mathcal{S}}\cap{\mathcal{C}}_{\mathcal{A}}$. To show equality it therefore suffices to show that an object of ${\mathcal{C}}_{\mathcal{A}}$ lies in ${{}^{\perp}}{\mathcal{F}}_{\mathcal{S}}$ if and only if it lies in ${\mathcal{W}}_{\mathcal{S}}$. This is proved via the usual argument using 2.16. The statements on the Quillen functors are proved dually to (i).
@@
Finally, ${\operatorname{id}}:{{}^{\ddagger}}{\mathcal{A}}^{i}_{\mathcal{S}}\to{{}^{\ddagger}}{\mathcal{A}}^{p}_{\mathcal{S}}$ will by definition be a right Quillen equivalence if, for $X\in{\mathcal{W}}_{\mathcal{S}}\cap{\mathcal{C}}_{\mathcal{A}}$ and $Y\in{\mathcal{W}}_{\mathcal{S}}\cap{\mathcal{F}}_{\mathcal{A}}$, a map $f:X\to Y$ is a weak equivalence in ${{}^{\ddagger}}{\mathcal{A}}^{p}_{\mathcal{S}}$ if and only if it is a weak equivalence in ${{}^{\ddagger}}{\mathcal{A}}^{i}_{\mathcal{S}}$. Note that the trivial cofibrations in ${{}^{\ddagger}}{\mathcal{A}}^{p}_{\mathcal{S}}$ are precisely the monomorphisms with projective cokernel, and so are still weak equivalences in ${{}^{\ddagger}}{\mathcal{A}}^{i}_{\mathcal{S}}$. Thus, by factorising $f$ in ${{}^{\ddagger}}{\mathcal{A}}^{p}_{\mathcal{S}}$ as a trivial cofibration followed by a fibration we may assume that $f$ is a fibration. Then it is a weak equivalence in ${{}^{\ddagger}}{\mathcal{A}}^{p}_{\mathcal{S}}$ if and only if its kernel lies in ${\mathcal{F}}_{\mathcal{S}}$. But since both $X$ and $Y$ lie in ${\mathcal{W}}_{\mathcal{S}}$ by assumption and ${\mathcal{W}}_{\mathcal{S}}$ is thick, this happens if and only if $\ker(f)\in{\mathcal{W}}_{\mathcal{S}}\cap{\mathcal{F}}_{\mathcal{S}}={\mathcal{W}}_{\mathcal{A}}$, cf. 4.5. Similarly, $f$ is a weak equivalence in ${{}^{\ddagger}}{\mathcal{A}}^{i}_{\mathcal{S}}$ if and only if $\ker(f)\in{\mathcal{W}}_{\mathcal{S}}\cap{\mathcal{C}}_{\mathcal{S}}={\mathcal{W}}_{\mathcal{A}}$. ∎
@@
In particular, by considering ${\mathcal{A}}={\operatorname{Mod}}({\mathcal{H}})$ with its Gorenstein projective and injective model structures and by letting =${\mathcal{M}}=\{M_{1},\ldots,M_{n}\}$ be some fixed collection of standard Levi subgroups of $G$, the above tells us that the recollement in 3.13(i) may be reconstructed from various model structures on ${\operatorname{Mod}}({\mathcal{H}})$ by only using the Gorenstein projective model structure. If we define
@@
where $F_{M}({\mathfrak{m}}):={\mathbb{X}}_{M}\otimes_{{\mathcal{H}}_{M}}{\mathfrak{m}}$ and $U_{M}(V):={\operatorname{Hom}}_{M}({\mathbb{X}}_{M},V)\cong V^{I_{M}}$ for any ${\mathfrak{m}}\in{\operatorname{Mod}}({\mathcal{H}}_{M})$ and $V\in{\operatorname{Rep}}_{k}^{\infty}(M)$. When $M=G$, we drop the subscript and write $F$ and $U$ for the corresponding functors. On the smooth representations side, the parabolic induction functor ${\operatorname{Ind}}_{P}^{G}:{\operatorname{Rep}}_{k}^{\infty}(M)\to{\operatorname{Rep}}_{k}^{\infty}(G)$ is defined by first inflating smooth representations of $M\cong P/N$ to $P$ and then applying the smooth induction functor. This functor has both a left and right adjoint (cf. [34, Proposition 4.3]), and its left adjoint can be described explicitly as the Jacquet functor $V\mapsto J_{P}^{G}(V):=V_{N}$ where $V_{N}$ denotes the $N$-coinvariants of $V$. This construction is compatible with the parabolic induction for ${\mathcal{H}}$-modules by a result of Ollivier-Vignéras.
Theorem 4.10 ([30, Theorem 1]).
@@
For any collection ${\mathcal{M}}$ of standard Levi subgroups of $G$, the right transfer of the model structures ${\operatorname{Mod}}({\mathcal{H}})^{G}_{\mathcal{M}}$ and ${{}^{\ddagger}}{\operatorname{Mod}}({\mathcal{H}})^{G}_{\mathcal{M}}$ to ${\operatorname{Rep}}_{k}^{\infty}(G)$ along $U$ both exist.